OpenStudy (anonymous):

Solve by elimination: 9x^2+y^2=85 2x^2-3y^2=6

OpenStudy (anonymous):

(3,2)?

OpenStudy (anonymous):

multiply the first equation by 3, and then add the second. 27x^2+3y^2=255 +2x^2-3y^2=6 29x^2=261 x^2=9 x=+3, -3. Back to the second equation 2(9)-3y^2=6 -3y^2=-12 y^2=4 y=2, -2. So our answers are (3,2), (3,-2), (-3,2), and (-3, -2)

OpenStudy (dumbcow):

yes but include the negative cases as well since you're dealing with x^2 and y^2 x = +-3 y = +-2 4 solutions (3,2) (3,-2) (-3,2) (-3,-2)

OpenStudy (anonymous):

Thanks guys :) Really appreciate the help